### v2

@@ Abstract
Keywords: Polychoric correlation, model misspecification, robust estimation, careless responding, rating scales
@@ 1 Introduction
Ordinal data are ubiquitous in psychology and related fields. With such data, e.g., arising from responses to rating scales, it is often recommended to estimate correlation matrices through polychoric correlation coefficients (e.g., Foldnes & Grønneberg,, 2022; Garrido et al.,, 2013; Holgado-Tello et al.,, 2010). The resulting polychoric correlation matrix is an important building block in subsequent multivariate models like factor analysis models and structural equation models (SEMs), as well as in explanatory methods like principal component analysis, multidimensional scaling, and clustering techniques (see, e.g., Mair,, 2018, for an overview). An individual polychoric correlation coefficient is the population correlation between two underlying latent variables that are postulated to have generated the observed categorical data through an unobserved discretization process. Traditionally, it is assumed that the two latent variables are standard bivariate normally distributed (Pearson,, 1901) to estimate the polychoric correlation coefficient from observed ordinal data. Estimation of this latent normality model, called the polychoric model, is commonly carried out via maximum likelihood (Olsson,, 1979). However, recent work has demonstrated that maximum likelihood (ML) estimation of polychoric correlation is highly sensitive to violations of the assumption of underlying normality. Violations of this assumption result in a misspecified polychoric model, which can lead to substantially biased estimates of its parameters and those of subsequent multivariate models, particularly SEMs (Grønneberg & Foldnes,, 2022; Foldnes & Grønneberg,, 2022, 2020, 2019).
Motivated by the recent interest in non-robustness of ML, we study estimation of the polychoric model under a misspecification framework stemming from the robust statistics literature (e.g., Huber & Ronchetti,, 2009). In this setup, which we call partial misspecification here, the polychoric model is potentially misspecified for an unknown (and possibly zero-valued) fraction of observations. Heuristically, the model is misspecified such that the affected subset of observations contains little to no relevant information for the parameter of interest, the polychoric correlation coefficient. Examples of such uninformative observations include careless responses, misresponses, or responses due to item misunderstanding. Especially careless responding has been identified as a major threat to the validity of questionnaire-based research findings (e.g., Huang et al., 2015b, ; Meade & Craig,, 2012; Credé,, 2010; Woods,, 2006). We demonstrate that already a small fraction of uninformative observations (such as careless respondents) can result in considerably biased ML estimates.
@@
The partial misspecification framework in this paper is fundamentally different to that considered in recent literature on misspecified polychoric models. In this literature (e.g., Grønneberg & Foldnes,, 2022; Foldnes & Grønneberg,, 2022, 2020, 2019), the polychoric model is misspecified in the sense that all (unobserved) realizations of the latent continuous variables come from a distribution that is nonnormal. Under this framework, which is also known as distributional misspecification, the parameter of interest is the correlation coefficient of the latent nonnormal distribution, and all observations are informative for this parameter. While the distributional misspecification framework led to novel insights regarding (the lack of) robustness in ML estimation of polychoric correlation, the partial misspecification framework of this paper can provide complimentary insights regarding the effects of a fraction of uninformative observations in a sample (such as careless responses), which is our primary objective.
@@ 2 Related literature
ML estimation of polychoric correlations was originally believed to be fairly robust to slight to moderate distributional misspecification (Li,, 2016; Coenders et al.,, 1997; Flora & Curran,, 2004; Maydeu-Olivares,, 2006). This belief was based on simulations that generated data for nonnormal latent variables via the Vale-Maurelli method (Vale & Maurelli,, 1983) that were then discretized to ordinal data. However, Grønneberg & Foldnes, (2019) show that the distribution of ordinal data generated in this way is indistinguishable from that of ordinal data stemming from discretizing normally distributed latent variables. In other words, simulation studies that ostensibly modeled nonnormality did in fact model normality. Simulating ordinal data in a way that ensures proper violations of latent normality (Grønneberg & Foldnes,, 2017) reveals that polychoric correlation is in fact highly susceptible to distributional misspecification, resulting in possibly large biases (Foldnes & Grønneberg,, 2020, 2022; Grønneberg & Foldnes,, 2022). Consequently, it is recommended to test for the validity of the latent normality assumption, for instance by using the bootstrap test of Foldnes & Grønneberg, (2020).
Another source of model misspecification occurs when the polychoric model is only misspecified for an uninformative subset of a sample (partial misspecification), where, in the context of this paper, the term “uninformative” refers to an absence of relevant information for polychoric correlation, for instance in careless responses. Careless responding “occurs when participants are not basing their response on the item content”, for instance when a participant is “unmotivated to think about what the item is asking” (Ward & Meade,, 2023). It has been shown to be a major threat to the validity of research results through a variety of psychometric issues, such as reduced scale reliability (Arias et al.,, 2020) and construct validity (Kam & Meyer,, 2015), attenuated factor loadings, improper factor structure, and deteriorated model fit in factor analyses (Arias et al.,, 2020; Huang et al., 2015a, ; Woods,, 2006), as well as inflated type I or type II errors in hypothesis testing (Arias et al.,, 2020; Huang et al., 2015b, ; Maniaci & Rogge,, 2014; McGrath et al.,, 2010; Woods,, 2006). Careless responding is widely prevalent (Ward & Meade,, 2023; Bowling et al.,, 2016; Meade & Craig,, 2012) with most estimates on its prevalence ranging from 10–15% of study participants (Curran,, 2016; Huang et al., 2015b, ; Huang et al.,, 2012; Meade & Craig,, 2012), while already a prevalence 5–10% can jeopardize the validity of research findings (Arias et al.,, 2020; Credé,, 2010; Woods,, 2006). In fact, Ward & Meade, (2023) conjecture that careless responding is likely present in all survey data. However, to the best of our knowledge, the effects of careless responding on estimates of the polychoric model have not yet been studied.
@@ 3.1 The polychoric model
$X=\begin{cases}1&\text{if }\xi<a_{1},\\ 2&\text{if }a_{1}\leq\xi<a_{2},\\ 3&\text{if }a_{2}\leq\xi<a_{3},\\ \vdots\\ K_{X}&\text{if }a_{K_{X}-1}\leq\xi,\\ \end{cases}$ and $Y=\begin{cases}1&\text{if }\eta<b_{1},\\ 2&\text{if }b_{1}\leq\eta<b_{2},\\ 3&\text{if }b_{2}\leq\eta<b_{3},\\ \vdots\\ K_{Y}&\text{if }b_{K_{Y}-1}\leq\eta,\\ \end{cases}$ (1)
@@
$\begin{pmatrix}\xi\\ \eta\end{pmatrix}\sim\textnormal{N}_{2}\Bigg{(}\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\Bigg{)}.$ (2)
@@
To distinguish arbitrary parameters $\bm{\theta}$ from a specific value under which the polychoric model generates ordinal data, denote the latter by $\bm{\theta}_{*}=\left(\rho_{*},a_{*,1},\dots,a_{*,K_{X}-1},b_{*,1},\dots,b_{*,K_{Y}-1}\right)^{\top}$. Given a random sample of ordinal data generated by a polychoric model under parameter value $\bm{\theta}_{*}$, the statistical problem is to estimate the true $\bm{\theta}_{*}$, which is traditionally achieved by the maximum likelihood estimator of Olsson, (1979).
@@ 3.2 Maximum likelihood estimation
$\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}=\arg\max_{\bm{\theta}\in\bm{\Theta}}\bigg{\{}\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}N_{xy}\log\left(p_{xy}(\bm{\theta})\right)\bigg{\}},$ (4)
@@ 4.3 Distributional misspecification
A model is distributionally misspecified when all observations in a given sample are generated by a distribution that is different from the model distribution. In the context of the polychoric model, this means that all ordinal observations are generated by a latent distribution that is nonnormal. Let $G$ denote the unknown nonnormal distribution that the latent variables $(\xi,\eta)$ jointly follow under distributional misspecification. The object of interest is the population correlation between latent $\xi$ and $\eta$ under distribution $G$, for which the normality-based ML estimator of Olsson, (1979) turns out to be substantially biased in many cases (e.g., Foldnes & Grønneberg,, 2020, 2022). As such, distributional misspecification is fundamentally different from partial misspecification: In the former, one attempts to estimate the population correlation of the nonnormal and unknown distribution that generated a sample, instead of estimating the polychoric correlation coefficient (which is the correlation under standard bivariate normality). In the latter, one attempts to estimate the polychoric correlation coefficient with a sample that has only been partly generated by latent normality (that is, the polychoric model). The assumption that the polychoric model is only partially misspecified for some uninformative observations enables one to still estimate the polychoric correlation coefficient of that model, which would not be feasible under distributional misspecification (at least not without additional assumptions).
@@ 5 Robust estimation of polychoric correlation
The consistency of ML estimation of an arbitrary model crucially depends on correct specification of that model. Indeed, ML estimation can be severely biased even when the assumed model is only slightly misspecified (e.g., Huber & Ronchetti,, 2009; Hampel et al.,, 1986; Huber,, 1964). For instance, in many models of continuous variables like regression models, one single observation from a different distribution can be enough to make the ML estimator converge to an arbitrary value (Huber & Ronchetti,, 2009, see also Alfons et al.,, 2022, for the special case of mediation analysis). The non-robustness of ML estimation of the polychoric model has been demonstrated empirically by Foldnes & Grønneberg, (2020, 2022); Grønneberg & Foldnes, (2022) for the case of distributional misspecification. In this section, we introduce an estimator that is designed to be robust to partial misspecification when present, but remains (asymptotically) equivalent to the ML estimator of Olsson, (1979) when misspecification is absent. We furthermore derive the statistical properties of the proposed estimator.
@@ 5.1 The estimator
as $N\to\infty$ (see, e.g., Chapter 19.2 in Van der Vaart,, 1998). If the polychoric model is correctly specified $(\varepsilon=0)$, then $\widehat{f}_{N}(x,y)$ will converge to the true model probability $p_{xy}(\bm{\theta}_{*})$ because
@@
is referred to as Pearson residual (PR, Lindsay,, 1994). It takes values in $[-1,+\infty)$ and can be interpreted as a goodness-of-fit measure. PR values close to 0 indicate a good fit between data and polychoric model at $\bm{\theta}$, whereas values toward $-1$ or +$\infty$+ ∞ indicate a poor fit because empirical response probabilities disagree with their model counterparts. To achieve robustness to misspecification of the polychoric model, responses that cannot be modeled well by the polychoric model, as indicated by their PR being away from 0, should receive less weight in the estimation procedure such that they do not over-proportionally affect the fit. Downweighting when necessary is achieved through a specific choice of discrepancy function $\varphi(\cdot)$ proposed by Welz, (2024), which is a special case of a function suggested by Ruckstuhl & Welsh, (2001). The discrepancy function reads
@@ 5.2 Statistical properties
We first address what the generalized ML estimator $\widehat{\bm{\theta}}_{N}$ in (10) actually estimates. Its estimand is given by
@@ 5.3 Implementation
An important user choice is that of the tuning constant $c$ in discrepancy function (11). The closer $c$ is to 0, the more robust the estimator will be to possible misspecification of the polychoric model (see, e.g., Figure 3). On the other hand, in the presence of model misspecification, the more robust the estimator is made, the larger its estimation variance becomes. Moreover, if the model is correctly specified, then Welz, (2024) shows that the most robust choice, $c=0$, is associated with two drawbacks, namely asymptotic nonnormality as well as certain finite sample issues. To avoid these drawbacks, we suggest to choose a value slightly larger than 0. In simulation experiments, we found that the choice $c=0.6$ yields a good compromise between robustness and efficiency in case of misspecification while not suffering from finite sample issues if the model is correctly specified. We therefore chose to use this value for all applications in this paper. However, we acknowledge that a detailed study, possibly a theoretical one, is necessary to provide guidelines on the choice of $c$. We will explore this in future work.
@@ 6 Simulation studies on partial misspecification
In this section, we employ two simulation studies to demonstrate the robustness gain of our proposed estimator under partial misspecification of the polychoric model. The first simulation design is intended to illustrate the performance of the robust estimator for the estimation of an individual polychoric correlation coefficient. The second design considers estimation of a polychoric correlation matrix, which is often am important building block in subsequent multivariate analyses, such a structural equation modeling or principal component analyses.

### v3

@@ Abstract
Keywords: Polychoric correlation, model misspecification, robust estimation, careless responding
@@ 1 Introduction
Ordinal data are ubiquitous in psychology and related fields. With such data, e.g., arising from responses to rating scales, it is often recommended to estimate correlation matrices through polychoric correlation coefficients (e.g., Foldnes & Grønneberg,, 2022; Garrido et al.,, 2013; Holgado-Tello et al.,, 2010). The resulting polychoric correlation matrix is an important building block in subsequent multivariate models like factor analysis models and structural equation models (SEMs), as well as in explanatory methods like principal component analysis, multidimensional scaling, and clustering techniques (see, e.g., Mair,, 2018, for an overview). An individual polychoric correlation coefficient is the population correlation between two underlying latent variables that are postulated to have generated the observed categorical data through an unobserved discretization process. Traditionally, it is assumed that the two latent variables are standard bivariate normally distributed (Pearson,, 1901) to estimate the polychoric correlation coefficient from observed ordinal data. Estimation of this latent normality model, called the polychoric model, is commonly carried out via maximum likelihood (Olsson,, 1979). However, recent work has demonstrated that maximum likelihood (ML) estimation of polychoric correlation is highly sensitive to violations of the assumption of underlying normality. Violations of this assumption result in a misspecified polychoric model, which can lead to substantially biased estimates of its parameters and those of subsequent multivariate models, particularly SEMs (Grønneberg & Foldnes,, 2022; Foldnes & Grønneberg,, 2022, 2020, 2019; Jin & Yang-Wallentin,, 2017).
Motivated by the recent interest in non-robustness of ML, we study estimation of the polychoric model under a misspecification framework stemming from the robust statistics literature (e.g., Huber & Ronchetti,, 2009). In this setup, which we call partial misspecification here, the polychoric model is potentially misspecified for an unknown (and possibly zero-valued) fraction of observations. Heuristically, the model is misspecified such that the affected subset of observations contains little to no relevant information for the parameter of interest, the polychoric correlation coefficient. Examples of such uninformative observations include careless responses, misresponses, or responses due to item misunderstanding. Especially careless responding has been identified as a major threat to the validity of questionnaire-based research findings (e.g., Welz et al.,, 2024; Huang et al., 2015b, ; Meade & Craig,, 2012; Credé,, 2010; Woods,, 2006). We demonstrate that already a small fraction of uninformative observations (such as careless respondents) can result in considerably biased ML estimates.
@@
The partial misspecification framework in this paper is fundamentally different to that considered in recent literature on misspecified polychoric models. In this literature (e.g., Lyhagen & Ornstein,, 2023; Grønneberg & Foldnes,, 2022; Foldnes & Grønneberg,, 2022, 2020, 2019; Jin & Yang-Wallentin,, 2017), the polychoric model is misspecified in the sense that all (unobserved) realizations of the latent continuous variables come from a distribution that is nonnormal. Under this framework, which is also known as distributional misspecification, the parameter of interest is the correlation coefficient of the latent nonnormal distribution, and all observations are informative for this parameter. While the distributional misspecification framework led to novel insights regarding (the lack of) robustness in ML estimation of polychoric correlation, the partial misspecification framework of this paper can provide complimentary insights regarding the effects of a fraction of uninformative observations in a sample (such as careless responses), which is our primary objective.
@@ 2 Related literature
ML estimation of polychoric correlations was originally believed to be fairly robust to slight to moderate distributional misspecification (Li,, 2016; Coenders et al.,, 1997; Flora & Curran,, 2004; Maydeu-Olivares,, 2006). This belief was based on simulations that generated data for nonnormal latent variables via the Vale-Maurelli method (Vale & Maurelli,, 1983) that were then discretized to ordinal data. However, Grønneberg & Foldnes, (2019) show that the distribution of ordinal data generated in this way is indistinguishable from that of ordinal data stemming from discretizing normally distributed latent variables. In other words, simulation studies that ostensibly modeled nonnormality did in fact model normality. Simulating ordinal data in a way that ensures proper violations of latent normality (Grønneberg & Foldnes,, 2017) reveals that polychoric correlation is in fact highly susceptible to distributional misspecification, resulting in possibly large biases (Foldnes & Grønneberg,, 2020, 2022; Grønneberg & Foldnes,, 2022; Jin & Yang-Wallentin,, 2017). Consequently, it is recommended to test for the validity of the latent normality assumption, for instance by using the bootstrap test of Foldnes & Grønneberg, (2020).
Another source of model misspecification occurs when the polychoric model is only misspecified for an uninformative subset of a sample (partial misspecification), where, in the context of this paper, the term “uninformative” refers to an absence of relevant information for polychoric correlation, for instance in careless responses. Careless responding “occurs when participants are not basing their response on the item content”, for instance when a participant is “unmotivated to think about what the item is asking” (Ward & Meade,, 2023). It has been shown to be a major threat to the validity of research results through a variety of psychometric issues, such as reduced scale reliability (Arias et al.,, 2020) and construct validity (Kam & Meyer,, 2015), attenuated factor loadings, improper factor structure, and deteriorated model fit in factor analyses (Arias et al.,, 2020; Huang et al., 2015a, ; Woods,, 2006), as well as inflated type I or type II errors in hypothesis testing (Arias et al.,, 2020; Huang et al., 2015b, ; Maniaci & Rogge,, 2014; McGrath et al.,, 2010; Woods,, 2006). Careless responding is widely prevalent (Ward & Meade,, 2023; Bowling et al.,, 2016; Meade & Craig,, 2012) with most estimates on its prevalence ranging from 10–15% of study participants (Curran,, 2016; Huang et al., 2015b, ; Huang et al.,, 2012; Meade & Craig,, 2012), while already a prevalence 5–10% can jeopardize the validity of research findings (Welz et al.,, 2024; Arias et al.,, 2020; Credé,, 2010; Woods,, 2006). In fact, Ward & Meade, (2023) conjecture that careless responding is likely present in all survey data. However, to the best of our knowledge, the effects of careless responding on estimates of the polychoric model have not yet been studied.
@@ 3.1 The polychoric model
$X=\begin{cases}1&\text{if }\xi<a_{1},\\ 2&\text{if }a_{1}\leq\xi<a_{2},\\ 3&\text{if }a_{2}\leq\xi<a_{3},\\ \vdots\\ K_{X}&\text{if }a_{K_{X}-1}\leq\xi,\\ \end{cases}\qquad\textnormal{ and }\qquad Y=\begin{cases}1&\text{if }\eta<b_{1},\\ 2&\text{if }b_{1}\leq\eta<b_{2},\\ 3&\text{if }b_{2}\leq\eta<b_{3},\\ \vdots\\ K_{Y}&\text{if }b_{K_{Y}-1}\leq\eta,\\ \end{cases}$ (1)
@@
$\begin{pmatrix}\xi\\ \eta\end{pmatrix}\sim\textnormal{N}_{2}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right).$ (2)
@@
To distinguish arbitrary parameters $\bm{\theta}$ from a specific value under which the polychoric model generates ordinal data, denote the latter by $\bm{\theta}_{*}=\left(\rho_{*},a_{*,1},\dots,a_{*,K_{X}-1},b_{*,1},\dots,b_{*,K_{Y}-1}\right)^{\top}$. Given a random sample of ordinal data generated by a polychoric model under parameter value $\bm{\theta}_{*}$, the statistical problem is to estimate the true $\bm{\theta}_{*}$, which is traditionally achieved by the maximum likelihood estimator of Olsson, (1979). [^1] Alternatives to the commonly used maximum likelihood estimator of Olsson, (1979) have been proposed in the literature, see, for instance, Zhang et al., (2024), Jöreskog, (1994), and Lau, (1985).
@@ 3.2 Maximum likelihood estimation
$\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}=\arg\max_{\bm{\theta}\in\bm{\Theta}}\left\{\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}N_{xy}\log\left(p_{xy}(\bm{\theta})\right)\right\},$ (4)
@@ 4.3 Distributional misspecification
A model is distributionally misspecified when all observations in a given sample are generated by a distribution that is different from the model distribution. In the context of the polychoric model, this means that all ordinal observations are generated by a latent distribution that is nonnormal. Let $G$ denote the unknown nonnormal distribution that the latent variables $(\xi,\eta)$ jointly follow under distributional misspecification. The object of interest is the population correlation between latent $\xi$ and $\eta$ under distribution $G$, for which the normality-based ML estimator of Olsson, (1979) turns out to be substantially biased in many cases (e.g., Lyhagen & Ornstein,, 2023; Foldnes & Grønneberg,, 2020, 2022; Jin & Yang-Wallentin,, 2017). As such, distributional misspecification is fundamentally different from partial misspecification: In the former, one attempts to estimate the population correlation of the nonnormal and unknown distribution that generated a sample, instead of estimating the polychoric correlation coefficient (which is the correlation under standard bivariate normality). In the latter, one attempts to estimate the polychoric correlation coefficient with a sample that has only been partly generated by latent normality (that is, the polychoric model). The assumption that the polychoric model is only partially misspecified for some uninformative observations enables one to still estimate the polychoric correlation coefficient of that model, which would not be feasible under distributional misspecification (at least not without additional assumptions).
@@ 5 Robust estimation of polychoric correlation
The consistency of ML estimation of an arbitrary model crucially depends on correct specification of that model. Indeed, ML estimation can be severely biased even when the assumed model is only slightly misspecified (e.g., Huber & Ronchetti,, 2009; Hampel et al.,, 1986; Huber,, 1964). For instance, in many models of continuous variables like regression models, one single observation from a different distribution can be enough to make the ML estimator converge to an arbitrary value (Huber & Ronchetti,, 2009); see also Alfons et al., (2022) for the special case of mediation analysis. The non-robustness of ML estimation of the polychoric model has been demonstrated empirically by Foldnes & Grønneberg, (2020, 2022); Grønneberg & Foldnes, (2022) for the case of distributional misspecification. In this section, we introduce an estimator that is designed to be robust to partial misspecification when present, but remains (asymptotically) equivalent to the ML estimator of Olsson, (1979) when misspecification is absent. We furthermore derive the statistical properties of the proposed estimator.
@@ 5.1 The estimator
as $N\to\infty$ (see, e.g., Chapter 19.2 in Van der Vaart,, 1998). If the polychoric model is correctly specified $(\varepsilon=0)$, then $\widehat{f}_{N}(x,y)$ will converge (in probability) to the true model probability $p_{xy}(\bm{\theta}_{*})$ because
@@
is referred to as Pearson residual (PR) (Lindsay,, 1994). It takes values in $[-1,+\infty)$ and can be interpreted as a goodness-of-fit measure. PR values close to 0 indicate a good fit between data and polychoric model at $\bm{\theta}$, whereas values toward $-1$ or +$\infty$+ ∞ indicate a poor fit because empirical response probabilities disagree with their model counterparts. To achieve robustness to misspecification of the polychoric model, responses that cannot be modeled well by the polychoric model, as indicated by their PR being away from 0, should receive less weight in the estimation procedure such that they do not over-proportionally affect the fit. Downweighting when necessary is achieved through a specific choice of discrepancy function $\varphi(\cdot)$ proposed by Welz, (2024), which is a special case of a function suggested by Ruckstuhl & Welsh, (2001). The discrepancy function reads
@@ 5.2 Statistical properties
We first address what the generalized ML estimator $\widehat{\bm{\theta}}_{N}$ in (10) estimates. Its estimand is given by
@@ 5.3 Implementation
An important user choice is that of the tuning constant $c$ in discrepancy function (11). The closer $c$ is to 0, the more robust the estimator will be to possible misspecification of the polychoric model (see, e.g., Figure 3). On the other hand, in the presence of model misspecification, the more robust the estimator is made, the larger its estimation variance becomes. Moreover, if the model is correctly specified, then Welz, (2024) shows that the most robust choice, $c=0$, is associated with two drawbacks, namely asymptotic nonnormality as well as certain finite sample issues. To avoid these drawbacks, we suggest to choose a value slightly larger than 0. In simulation experiments, we found that the choice $c=0.6$ yields a good compromise between robustness and efficiency in case of misspecification while not suffering from finite sample issues if the model is correctly specified. We therefore chose to use this value for all applications in this paper. However, we acknowledge that a detailed study, preferably founded in statistical theory, is necessary to provide guidelines on the choice of $c$. We will explore this in future work.
@@ 6 Simulation studies on partial misspecification
In this section, we employ two simulation studies to demonstrate the robustness gain of our proposed estimator under partial misspecification of the polychoric model. The first simulation design is a simplified setting with respect to the partial misspecification, chosen specifically to illustrate the effects of a particular misspecification type with high leverage affecting only a small number of contingency table cells. The second design is more involved and considers estimation of a polychoric correlation matrix with a misspecification type that scatters in many directions so that nonnormal data points can occur in every contingency table cell.